OpenStudy (anonymous):

i need help graphing y=x+7

jimthompson5910 (jim_thompson5910):

When x=0, then y=0+7 = 7 So we have one point (0,7) When x=1, then y=1+7 = 8 So we have one point (1,8) So plot the two points (0,7) and (1,8) and draw a straight line through them to graph y = x+7

OpenStudy (anonymous):

Thank you